# Evaluating Arabic Large Language Models: A Survey of Benchmarks, Methods, and Gaps

### v1

@@ 1 Introduction
In recent years, Large Language Models (LLMs) have demonstrated remarkable advances in natural language understanding and logical reasoning capabilities, bringing us closer to the vision of AGI Naveed et al. (2025). From the beginning of the transformer era, LLMs have shown potential beyond English, with numerous successes across multiple languages Huang et al. (2024b). The Arabic language, spoken by almost 500 million people globally [^1] World Bank (2024). Population, Total – Arab World. https://data.worldbank.org/indicator/SP.POP.TOTL?locations=1A, exemplifies the critical need for multilingual LLM capabilities. Consequently, both bilingual and Arabic-specific LLMs have been released as closed and open-source models Al-Khalifa et al. (2025). A critical component of LLM development is the availability of robust benchmarks that enable systematic evaluation of performance across tasks and domains. In parallel with progress in Arabic LLMs, a corresponding effort has emerged focused on creating comprehensive benchmarks for Arabic evaluation.
@@ 2.1 Arabic LLMs
State-of-the-art Arabic LLMs can be categorized into three types. Native models are trained from scratch exclusively on Arabic data, enabling Arabic-only interaction. Examples include Jais Sengupta et al. (2023), ArabianGPT Koubaa et al. (2024), and AraGPT Antoun et al. (2021). Multilingual models support multiple languages including Arabic, such as Qwen3 Yang et al. (2025), Gemma3 Team et al. (2025b), and Llama Grattafiori et al. (2024), as well as closed-source models like ChatGPT OpenAI (2023) and Claude Anthropic (2024). Adapted Arabic models apply continued pretraining or supervised fine-tuning to existing multilingual models to enhance Arabic performance, exemplified by AceGPT Huang et al. (2024a), SILMA Team (2024), Fanar Team et al. (2025a), and Falcon-Arabic TII (2025).
@@ 4.2 Domain Knowledge
Poetry and Linguistics. Fann or Flop Alghallabi et al. (2025) assesses poetry understanding through 6,984 poem-explanation pairs, evaluating metaphorical and figurative comprehension. Evaluation employs BLEU, chrF(++), BERTScore, and mDeBERTaV3 for character-level overlap and semantic alignment, with GPT-4o as judge for faithfulness, grammatical correctness, and interpretive depth.
@@ 5.1 Comprehensive Multi-Task Benchmarks
LAraBench Abdelali et al. (2024) pioneered systematic LLM benchmarking for Arabic NLP and speech processing, comparing task-specific models against general-purpose LLMs across 33 tasks spanning 61 datasets. It integrates earlier resources like ARCD and dialect identification corpora, bridging pre-LLM and LLM evaluation eras. Evaluation of GPT-3.5-turbo, GPT-4 OpenAI et al. (2024), BLOOMZ Muennighoff et al. (2022), Jais-13b-chat Sengupta et al. (2023) and Whisper Radford et al. (2023) under zero-shot and few-shot settings reveals that specialized models outperform LLMs in zero-shot scenarios, though larger LLMs with few-shot adaptation substantially narrow this gap.
BALSAM Al-Matham et al. (2025a) provides community-driven comprehensive evaluation emphasizing instruction-following across 78 tasks and 14 categories, totaling 52K examples covering summarization, question answering, information extraction, translation, classification, creative writing, and reasoning. It addresses data contamination through blind test sets and provides an integrated leaderboard. Notably, LLM-as-judge correlates more strongly with human judgments (0.824 - 0.977) than traditional metrics (BLEU, ROUGE, BERTScore). Datasets combine naturally authored, translated, prompted, and synthetic sources from xP3 Muennighoff et al. (2023), PromptSource Bach et al. (2022), SuperNaturalInstructions Wang et al. (2022), TruthfulQA Lin et al. (2022), and 16 newly developed datasets (1,755 prompts). Results show large closed-source models (GPT-4o, Gemini 2.0, DeepSeek V3) outperform smaller Arabic-centric models, with performance depending on tokenization quality, training data volume, and Arabic-specific fine-tuning. Limitations include potential cultural misalignment in translated datasets and insufficient assessment of multi-turn dialogue and hallucination.
@@ 5.2 Task-Specific Benchmarks
ORCA Elmadany et al. (2023) focuses on natural language understanding across Arabic’s linguistic diversity, consolidating 60 datasets into seven task clusters: sentence classification, structured prediction (NER, POS), semantic textual similarity, paraphrase identification, natural language inference, word sense disambiguation, and question answering. Evaluation employs task-appropriate metrics (accuracy, F1-score, Pearson correlation, exact match) across Modern Standard Arabic and dialects, comparing 18 multilingual and Arabic-specific pre-trained models. Its public leaderboard with rich metadata promotes transparency and reproducibility, providing a critical reference for understanding evolution toward modern LLM evaluation.
@@ 6 Culture and Dialects
Evaluation of cultural and dialectal understanding in Arabic LLMs has developed progressively, with each benchmark addressing gaps left by its predecessors. Jawaher Magdy et al. (2025) provided one of the earliest culturally grounded resources by compiling 10,037 Arabic proverbs drawn from online resources. Each proverb was annotated with its dialectal origin and idiomatic explanation, and cast into a QA format, enabling evaluation of figurative and cultural reasoning in dialectal Arabic a dimension overlooked in prior benchmarks. However, Jawaher was limited to proverbial knowledge. PALM Alwajih et al. (2025b) addressed this gap by curating a broader culturally inclusive dataset through a year-long community-driven effort spanning all 22 Arab countries. Covering prose, dialogue, and cultural expressions across 20 topics, it expanded evaluation beyond proverbs and tested inclusivity using perplexity and generation quality. Following the release of Palm, PalmX Alwajih et al. (2025a) released a benchmark focused on assessing the deep understanding of Arabic and Islamic culture. It contained two subtasks of MCQs in MSA covering traditions, food, history, religious practices, and expressions. Despite these advances, benchmarks had yet to capture Arab-specific commonsense reasoning. Commonsense Reasoning in Arab Culture Sadallah et al. (2025) filled this gap by introducing questions requiring culturally aware inference, complementing the focus on linguistic diversity in earlier work.
@@ 7 Target-Specific Tasks
Instruction-Following. CamelEval Qian et al. (2024) evaluates conversational abilities and instruction-following through 1,610 generative questions: 805 human-validated translations from AlpacaEval Dubois et al. (2023) and 805 synthetically generated using GPT-4 from culturally grounded textbooks. Evaluation employs LLM-as-judge computing win rates for open-ended generation.
@@ 8.3 Arabic LLM Leaderboards
The Open Arabic LLM Leaderboard (OALL) El Filali et al. (2024a) pioneered open-source rankings, initially using Alghafa, EXAMS_ar, and MMLU_ar. OALL v2 El Filali et al. (2025) transitioned to native benchmarks (ArabicMMLU, ALRAGE, AraTrust, MadinahQA), reflecting community consensus against translated content. BALSAM Al-Matham et al. (2025b) offers comprehensive evaluation across 78 tasks (52K samples) with private test sets preventing contamination, including both closed and open-source models. ILMAAM Nacar et al. (2025) specializes in culturally aligned evaluation using refined ArabicMMLU by ensuring religious sensitivity and social norms. The AraGen benchmark El Filali et al. (2024b) and its associated leaderboard adopt a 3C3H evaluation metric (Correctness, Completeness, Conciseness, Helpfulness, Honesty, Harmlessness), using LLM-as-judge, and combine it with a dynamic, blind-testing approaches to push for robust and fair benchmarking of Arabic LLMs.
@@ Appendix B Arabic NLP Benchmarks Timeline
2025 BALSAM Al-Matham et al. (2025a) MT • Transliteration • DialectMT • Simplification • QRewrite • Paraphrase • Intent • GrammarCorr • GenderRewrite • TextClass • Sentiment • Sarcasm • DialectID • Command • Summarization • SubjectGen • AnswerExt • SeqTag 26k 1.7k 24k 52k LLM-as-Judge • BLEU • ROUGE BENCH Partial Yes
@@
2025 PALM Alwajih et al. (2025b) Prose • Dialogue • Cultural Expressions 17.4k - - 17.4k LLM-as-Judge GEN Yes Yes
2025 PalmX Alwajih et al. (2025a) Shared Task • Arabic Culture MCQs • Islamic Culture MCQs - 6.4k - 6.4k Accuracy, Log-prob MCQ Yes No

### v2

@@ 1 Introduction
In recent years, Large Language Models (LLMs) have demonstrated remarkable advances in natural language understanding and logical reasoning capabilities, bringing us closer to the vision of AGI Naveed et al. (2025). From the beginning of the transformer era, LLMs have shown potential beyond English, with numerous successes across multiple languages Huang et al. (2025). The Arabic language, spoken by almost 500 million people globally [^1] World Bank (2024). Population, Total – Arab World. https://data.worldbank.org/indicator/SP.POP.TOTL?locations=1A, exemplifies the critical need for multilingual LLM capabilities. Consequently, both bilingual and Arabic-specific LLMs have been released as closed and open-source models Al-Khalifa et al. (2025). A critical component of LLM development is the availability of robust benchmarks that enable systematic evaluation of performance across tasks and domains. In parallel with progress in Arabic LLMs, a corresponding effort has emerged focused on creating comprehensive benchmarks for Arabic evaluation.
@@ 2.1 Arabic LLMs
State-of-the-art Arabic LLMs can be categorized into three types. Native models are trained from scratch exclusively on Arabic data, enabling Arabic-only interaction. Examples include Jais Sengupta et al. (2023), ArabianGPT Koubaa et al. (2024), and AraGPT Antoun et al. (2021). Multilingual models support multiple languages including Arabic, such as Qwen3 Yang et al. (2025), Gemma3 Team et al. (2025b), and Llama Grattafiori et al. (2024), as well as closed-source models like ChatGPT OpenAI (2023) and Claude Anthropic (2024). Adapted Arabic models apply continued pretraining or supervised fine-tuning to existing multilingual models to enhance Arabic performance, exemplified by AceGPT Huang et al. (2024), SILMA Team (2024), Fanar Team et al. (2025a), and Falcon-Arabic TII (2025).
@@ 4.2 Domain Knowledge
Poetry and Linguistics. Fann or Flop Alghallabi et al. (2025) assesses poetry understanding through 6,984 poem-explanation pairs, evaluating metaphorical and figurative comprehension. Evaluation uses BLEU, chrF(++), BERTScore, and mDeBERTaV3 for character-level overlap and semantic alignment, with GPT-4o as judge for faithfulness, grammatical correctness, and interpretive depth.
@@ 5.1 Comprehensive Multi-Task Benchmarks
LAraBench Abdelali et al. (2024) pioneered systematic LLM benchmarking for Arabic NLP and speech processing, comparing task-specific models against general-purpose LLMs across 33 tasks spanning 61 datasets. It integrates earlier resources like ARCD and dialect identification corpora, bridging pre-LLM and LLM evaluation eras. Evaluation of GPT-3.5-turbo, GPT-4 OpenAI et al. (2024), BLOOMZ Muennighoff et al. (2023a), Jais-13b-chat Sengupta et al. (2023) and Whisper Radford et al. (2023) under zero-shot and few-shot settings reveals that specialized models outperform LLMs in zero-shot scenarios, though larger LLMs with few-shot adaptation substantially narrow this gap.
BALSAM Al-Matham et al. (2025) provides community-driven comprehensive evaluation emphasizing instruction-following across 78 tasks and 14 categories, totaling 52K examples covering summarization, question answering, information extraction, translation, classification, creative writing, and reasoning. It addresses data contamination through blind test sets and provides an integrated leaderboard. Notably, LLM-as-judge correlates more strongly with human judgments (0.824 - 0.977) than traditional metrics (BLEU, ROUGE, BERTScore). Datasets combine naturally authored, translated, prompted, and synthetic sources from xP3 Muennighoff et al. (2023b), PromptSource Bach et al. (2022), SuperNaturalInstructions Wang et al. (2022), TruthfulQA Lin et al. (2022), and 16 newly developed datasets (1,755 prompts). Results show large closed-source models (GPT-4o, Gemini 2.0, DeepSeek V3) outperform smaller Arabic-centric models, with performance depending on tokenization quality, training data volume, and Arabic-specific fine-tuning. Limitations include potential cultural misalignment in translated datasets and insufficient assessment of multi-turn dialogue and hallucination.
@@ 5.2 Task-Specific Benchmarks
ORCA Elmadany et al. (2023) focuses on natural language understanding across Arabic’s linguistic diversity, consolidating 60 datasets into 7 task clusters: sentence classification, structured prediction (NER, POS), semantic textual similarity, paraphrase identification, natural language inference, word sense disambiguation, and question answering. Evaluation employs task-appropriate metrics (accuracy, F1-score, Pearson correlation, exact match) across MSA and dialects, comparing 18 multilingual and Arabic-specific pre-trained models. Its public leaderboard with rich metadata promotes transparency and reproducibility, providing a critical reference for understanding evolution toward modern LLM evaluation.
@@ 6 Culture and Dialects
Evaluation of cultural and dialectal understanding in Arabic LLMs has developed progressively, with each benchmark addressing gaps left by its predecessors. Jawaher Magdy et al. (2025) provided one of the earliest culturally grounded resources by compiling 10,037 Arabic proverbs drawn from online resources. Each proverb was annotated with its dialectal origin and idiomatic explanation, and cast into a QA format, enabling evaluation of figurative and cultural reasoning in dialectal Arabic a dimension overlooked in prior benchmarks. However, Jawaher was limited to proverbial knowledge.
PALM Alwajih et al. (2025a) addressed this gap by curating a broader culturally inclusive dataset through a year-long community-driven effort spanning all 22 Arab countries. Covering prose, dialogue, and cultural expressions across 20 topics, it expanded evaluation beyond proverbs and tested inclusivity using perplexity and generation quality. Following the release of Palm, PalmX Alwajih et al. (2025b) released a benchmark focused on assessing the deep understanding of Arabic and Islamic culture. It contained two subtasks of MCQs in MSA covering traditions, food, history, religious practices, and expressions. Despite these advances, benchmarks had yet to capture Arab-specific commonsense reasoning. Commonsense Reasoning in Arab Culture Sadallah et al. (2025) filled this gap by introducing questions requiring culturally aware inference, complementing the focus on linguistic diversity in earlier work.
@@ 7 Target-Specific Tasks
Instruction-Following. CamelEval Qian et al. (2024) evaluates conversational abilities and instruction-following through 1,610 generative questions: 805 human-validated translations from AlpacaEval Dubois et al. (2023) and 805 synthetically generated using GPT-4 from culturally grounded textbooks. Evaluation uses LLM-as-judge computing win rates for open-ended generation.
@@ 8.3 Arabic LLM Leaderboards
The Open Arabic LLM Leaderboard (OALL) El Filali et al. (2024a) pioneered open-source rankings, initially using Alghafa, EXAMS_ar, and MMLU_ar. OALL v2 El Filali et al. (2025) transitioned to native benchmarks (ArabicMMLU, ALRAGE, AraTrust, MadinahQA), reflecting community consensus against translated content. BALSAM Al-Matham et al. (2025) offers comprehensive evaluation across 78 tasks (52K samples) with private test sets preventing contamination, including both closed and open-source models. ILMAAM Nacar et al. (2025) specializes in culturally aligned evaluation using refined ArabicMMLU by ensuring religious sensitivity and social norms. The AraGen benchmark El Filali et al. (2024b) and its associated leaderboard adopt a 3C3H evaluation metric (Correctness, Completeness, Conciseness, Helpfulness, Honesty, Harmlessness), using LLM-as-judge, and combine it with a dynamic, blind-testing approaches to push for robust and fair benchmarking of Arabic LLMs.
@@ Appendix B Arabic NLP Benchmarks Timeline
2025 BALSAM Al-Matham et al. (2025) MT • Transliteration • DialectMT • Simplification • QRewrite • Paraphrase • Intent • GrammarCorr • GenderRewrite • TextClass • Sentiment • Sarcasm • DialectID • Command • Summarization • SubjectGen • AnswerExt • SeqTag 26k 1.7k 24k 52k LLM-as-Judge • BLEU • ROUGE BENCH Partial Yes
@@
2025 PALM Alwajih et al. (2025a) Prose • Dialogue • Cultural Expressions 17.4k - - 17.4k LLM-as-Judge GEN Yes Yes
2025 PalmX Alwajih et al. (2025b) Shared Task • Arabic Culture MCQs • Islamic Culture MCQs - 6.4k - 6.4k Accuracy, Log-prob MCQ Yes No